14 of 15 (0 complete) Evaluate the definite integral J (1x² - 6x + 5) dx (7x² - 6x + 5) dx = 0 (Simplify your answer.)
Answer #1

LO < =41549-05 : S = that I = 8+ € = s+&+= st ty+ & (0-135+–7), +($-f]l. [x] 5+ (4]o-lo] piss+upajo-v=f p/st*9=49) Š